OpenStudy (anonymous):

Solve exactly for the variable: \[x-\frac{ 1 }{ 2 }x^3+\frac{ 1 }{ 3 }x^3+...=0.2\] No idea how to do this. We are in taylor polynomial at the moment.

OpenStudy (anonymous):

the 2nd term should be x^2 \[x-\frac{ 1 }{ 2 }x^2+\frac{ 1 }{ 3 }x^3+...=0.2\]

OpenStudy (anonymous):

Any hints for starting this?